A current of 1 A exists in a copper wire of cross section 1 mm^2 . Assuming one free electron per atom , the drift speed of free electron ?

Grade:10

1 Answers

Vikas TU
6 years ago
As we realize that the course of float speed of conduction electrons is inverse to the electric field heading, 
i.e., electrons float toward expanding potential. The float speed is given by 
float speed =( I/neA)
Presently  e=1.6 \times 10^{-19}C,A= 1.0 \times 10^{-7}m^2, I =1.5A
The thickness of conduction electrons, $n$ is equivalent to the quantity of particles per cubic meter 
(accepting one conduction electron for every $ Cu$ iota as is sensible from its valence electron number of one). 
A cubic meter of copper has a mass of $9.0 \times 10^3 \: kg$. Since $6.0 \times 10^{23}$ copper particles have a mass of $ 63.5 g,
n = \large\frac{6 \times 10^{23}}{63.5} \times 9 \times 10^{6} = 8.5 \times 10^{28}
on substituting in the recipe for float speed ,one finds the last solution as 1.1 mm .